I am supposed to solve:
sqrt with the radicand 5x +21, then = x+3. So the question is:
sqrt 5x+21=x+3

Found 2 solutions by checkley77, stanbon:
Answer by checkley77(12844)

sqrt(5x+21)=x+3 square both sides.
5x+21=x^2+6x+9
x^2+6x-5x+9-21=0
x^2+x-12=0
(x+4)(x-3)=0
x+4=0
x=-4 ans.
x-3=0
x=3 ans.

Answer by stanbon(75887)

I am supposed to solve:
sqrt with the radicand 5x +21, then = x+3. So the question is:
sqrt(5x+21) = x+3
---
Square both sides to get:
5x+21 = x^2 + 6x + 9
---
Then x^2 + x - 12 = 0
(x+4)(x-3) = 0
x = -4 or x = 3
-----------------------
Since you squared, you have to check for extraneous answers.
------------------------------
Checking x = -4 you get:
sqrt(-20+21) = -4+3
sqrt(1) = -1
1 = -1
False ; so x = -4 is extraneous
-------------------------------------
Checking x = 3 you get:
sqrt(15+21) = 3+3
sqrt(36) = 6
6 = 6
-----------------------------
Final Answer: x = 3
===============================
Cheers,
Stan H.